how can you fill in a table in with y=x²-4
and also how do you graph it

Answer by Edwin McCravy(20055)

We start with this:

 x   Substitute    work      y    point to plot (x,y)  
-----------------------------------------------------
-4   y=(-4)²-4  =  16-4  =  12        (-4,12)
-3   y=(-3)²-4  =   9-4  =   5         (-3,5)
-2   y=(-2)²-4  =   4-4  =   0         (-2,0)
-1   y=(-1)²-4  =   1-4  =  -3        (-1,-3)
 0    y=(0)²-4  =   0-4  =  -4         (0,-4)
 1    y=(1)²-4  =   1-4  =  -3         (1,-3)
 2    y=(2)²-4  =   4-4  =   0          (2,0)
 3    y=(3)²-4  =   9-4  =   5          (3,5)
 4    y=(4)²-4  =  16-4  =  12         (4,12)  

Now we plot all those points:



Now we draw a smooth u-shaped graph through those points:



That's it!

Edwin